# Einstein manifolds of negative lower bounds on curvature operator of the second Kind

### v1

@@ 1. Introduction and main results
In this paper, we explore the sphere theorems for Einstein manifolds within the context of conditions imposed by the curvature operator of the second kind. The study of the curvature operator of the second kind, as defined in Section 2.2 below, can be traced back to a conjecture by Nishikawa [Nis86], which asserts that a closed Riemannian manifold with positive (or nonnegative) curvature operator of the second kind is diffeomorphic to a spherical space form (or a Riemannian locally symmetric space). Recently, Cao-Gursky-Tran [CGT23] and Li [Li24a] have confirmed Nishikawa’s conjecture for the positive and nonnegative cases, respectively. For further information and recent developments on this topic, we direct readers to the following references:[Li23a, Li22b, Li23b, Li24b, FL24, NPW23, NPWW23, DF24, DFY24].
@@
A closed Einstein manifold with nonnegative curvature operator of the second kind is either flat or a round sphere.
[…]
Cao-Gursky-Tran [CGT23], Nienhaus-Petersen-Wink [NPW23], Li [Li24a] and Dai-Fu [DF24] have proved many excellent results under various curvature conditions. Recently, Li [Li24b] introduced new lower bound conditions on the curvature operator of the second kind (see [Li24b, Definition 1.2]) and proved optimal sphere theorems. Based on this framework, we focus on Li’s condition with $\alpha=1$, namely (1.1), which is weaker than nonnegative curvature operator of the second kind. To state our result, we introduce some notations: $(M^{n},g)$ is an $n$ dimensional Riemannian manifold, $\mathring{R}$ represents the curvature operator of the second kind, $N=(n+2)(n-1)/2$ is the dimension of $S_{0}^{2}(T_{p}M)$, $\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{N}$ are the eigenvalues of $\mathring{R}$, and $\bar{\lambda}=\sum_{j=1}^{N}\lambda_{j}/N$ is the average of $\lambda_{j}$. The main result of this paper is as follows:
@@
(1.1) $\lambda_{1}\geq-\theta(n)\bar{\lambda}.$
@@
(1.2) $\begin{split}\theta(n)=\frac{3(N-1)(N+1-n)}{(N-1)(N-3)+3n(N-2)}-1.\end{split}$
Remark 1.1.
$\theta(n)$ defined by (1.2) is strictly positive and increasing in $n\in[4,\infty)$ and tends to 2 as $n\rightarrow\infty$.
@@
The crucial step in proving Theorem 1.2 is solving an extremum problem introduced in Proposition 3.4. Upon examination of the proof, it becomes evident that for any $\theta$, strictly greater than $\theta(n)$, Proposition 3.4 does not hold. Consequently, by employing our method, $\theta(n)$ as provided in equation (1.2) is deemed optimal. But it remains an interesting question that whether $\theta(n)$ is sharp or not for Theorem 1.2.
@@ 2.1. Notation and Conventions
In this section, we set up some notations and recall some basic facts about the curvature operator of the second kind, and refer to [CV60, BK78, OT79, Nis86, Kas93, Li24a] for more details.
@@ 2.2. Curvature Operator of the Second Kind
For an algebraic curvature tensor $R\in S^{2}_{B}(\wedge^{2}(T_{p}M))$, $R$ can induce two self-adjoint operators (see [Nis86]). The first one denoted by $\hat{R}:\wedge^{2}(T_{p}M)\to\wedge^{2}(T_{p}M)$ is defined as
@@
The curvature operator of the second kind, introduced by Nishikawa [Nis86], is symmetric bilinear form
@@
obtained by restricting $\overline{R}$ to $S^{2}_{0}(T_{p}M)$, the space of traceless symmetric two-tensors. It was pointed out in [NPW23] that the curvature operator of the second kind $\mathring{R}$ can also be interpreted as the self-adjoint operator
@@
Suppose $(M^{n},g)$ is an Einstein manifold of dimension $n$. Then
@@ 2.3. A formula for Weyl tensor on Einstein manifolds
Given $A,B\in S^{2}(V)$, their Kulkarni-Nomizu product gives rise to $A\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}B\in S^{2}_{B}(\wedge^{2}V)$ via
$(A\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}B)_{ijkl}=A_{ik}B_{jl}+A_{jl}B_{ik}-A_{jk}B_{il}-A_{il}B_{jk}.$
It is well known that the Riemann curvature tensor $R$ can be decomposed into irreducible components (cf. [CM20, (1.79)]) as
(2.19) $R=W+\frac{1}{n-2}\operatorname{Ric}\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}g-\frac{{\operatorname{Scal}}}{2(n-1)(n-2)}g\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}g,$
@@
(2.28) $R=W+\frac{{\operatorname{Scal}}}{2n(n-2)}g\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}g.$
@@
We recall some curvature notations on $(M^{n},g)$: $R$ represents the Riemann curvature operator, $W$ represents the Weyl tensor, $\mathring{R}$ represents the curvature operator of the second kind, $\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{N}$ are the eigenvalues of $\mathring{R}$, and $\bar{\lambda}=\sum_{j=1}^{N}\lambda_{j}/N$ is the average of $\lambda_{j}$ To prove the main theorem, we conduct several computations involving the Riemann curvature through the curvature operator of the second kind on Einstein manifolds.
@@
Let $(M^{n},g)$ be an $n$ dimensional Einstein manifold. Then
@@
$R=\frac{{\operatorname{Scal}}}{2n(n-1)}g\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}g+W,$
@@
$|R|^{2}=\frac{{\operatorname{Scal}}^{2}}{4n^{2}(n-1)^{2}}|g\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}g|^{2}+|W|^{2}=\frac{{\operatorname{Scal}}^{2}}{2n(n-1)}+|W|^{2},$
where we used the identity $|g\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}g|^{2}=8n(n-1)$ (cf. [Lee18, Lemma 7.22]). Combining this with identity (2.2), we conclude
@@
Let $(M^{n},g)$ be an $n$-dimensional Einstein manifold. Suppose the smallest eigenvalue of $\mathring{R}$ satisfies $\lambda_{1}\geq-\theta\bar{\lambda}$, then
@@
and by the assumption $\lambda_{1}\geq-\theta\bar{\lambda}$, we estimate that
@@
Suppose $(M^{n},g)$ is an Einstein manifold satisfying $\lambda_{1}\geq-\theta\bar{\lambda}$. Then
@@
=$\quad$= $\frac{16}{3n}\Big{[}N(N-3)\theta-(2N-9n+6)N\Big{]}\bar{\lambda}^{3}$
@@
Let $\theta$ be the positive constant defined by (1.2). If $\bar{\lambda}>0$ and $-\theta\bar{\lambda}\leq\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{N}$, then
$f(\lambda)\geq 0,$
and the equality holds if and only if
$\lambda=(1,1,\cdots,1)\bar{\lambda}\text{\quad or \quad}\lambda=(-\theta,\frac{N+\theta}{N-1},\cdots,\frac{N+\theta}{N-1})\bar{\lambda}.$
@@
From the expression of $f(\lambda)$ we can deduce that
@@
$\sum_{j=1}^{N}x_{j}=Nt,\quad\text{and}\quad x_{j}\geq 0\ (j=1,2,\cdots N).$
@@
When $l$ equals $\frac{k}{2}$ and equation (3.7) is solvable, we have $a+b=\frac{2N}{k}t$ and $1-\frac{N-2}{3(N-1)}=\frac{N}{k}$, and in this case direct calculation reveals that
(3.8) $F\big{(}{\underbrace{0,\cdots,0}_{N-k},\underbrace{a,\cdots,a}_{k/2},\underbrace{b,\cdots,b}_{k/2}}\big{)}=F(P_{k,0})$
for $\forall a,b\geq 0$ satisfying $a+b=2Nt/k$.
When $l<\frac{k}{2}$, by solving (3.7) we derive
@@
Therefore we conclude from (3.8) and (3.9) that
$\min\Big{\{}F(x):\sum_{j=1}^{n}x_{j}=Nt,x_{j}\geq 0\Big{\}}=\min\Big{\{}F(P_{k,0}),k=1,2,\cdots,N\Big{\}},$
@@
$\lambda^{m}=(\underbrace{-\theta\bar{\lambda},\cdots,-\theta\bar{\lambda}}_{m},\underbrace{\frac{N+m\theta}{N-m}\bar{\lambda},\cdots,\frac{N+m\theta}{N-m}\bar{\lambda}}_{N-m})$
@@
From the definition of $\theta$ in (1.2), we see
@@
and substituting above to (3.10), we find
$f(\lambda^{0})=f(\lambda^{1})=0,\quad\text{and}\quad f(\lambda^{k})>f(\lambda^{1})=0\quad(k=2,3,\cdots,N-1).$
Hence Proposition 3.4 comes true. $\square$
Proof of the Theorem 1.2.
If $\bar{\lambda}>0$, then Lemma 3.3 and Proposition 3.4 imply
(3.11) $\langle\Delta R,R\rangle\geq 0.$
If $\bar{\lambda}=0$, $\lambda_{1}\geq-\theta(n)\bar{\lambda}$ implies $\lambda_{1}=\lambda_{2}=\cdots\lambda_{N}=0$, and (3.11) holds by Lemma 3.3. Using (3.11), we estimate that
@@
and integrating above inequality over $M$ yields
[…]
which implies $|\nabla R|=0$. Therefore $M$ is a symmetric space. [Li24b, Proposition 2.8] and [Li24b, Theorem 1.8(2)] imply that $M$ is either flat or a rational homology sphere. Besides, Poincaré duality shows that rational cohomology is determined by rational homology [Hat02, Chapter 3.3]. Thus we conclude that $M$ is either flat or a symmetric space which is also a rational cohomology sphere. Compact symmetric spaces which are real cohomology spheres were classified by Wolf in [Wol69, Theorem 1]. If $\bar{\lambda}>0$, apart from spheres, $SU(3)/SO(3)$ is the only simply connected example. Proposition 3.4 implies that the eigenvalues of $\mathring{R}$ is either $\lambda=(1,\cdots,1)\bar{\lambda}$ or $\lambda=(-\theta,\frac{N+\theta}{N-1},\cdots,\frac{N+\theta}{N-1})\bar{\lambda}$, and combining with [NPW23, Example 4.5], we know that $M$ is a round sphere. $\square$

### v2

@@ 1. Introduction and main results
In this paper, we explore the sphere theorems for Einstein manifolds within the context of conditions imposed by the curvature operator of the second kind. The study of the curvature operator of the second kind, as defined in Section 2.2 below, can be traced back to a conjecture by Nishikawa [Nishikawa86], which asserts that a closed Riemannian manifold with positive (or nonnegative) curvature operator of the second kind is diffeomorphic to a spherical space form (or a Riemannian locally symmetric space). Recently, Cao-Gursky-Tran [CGT23] and Li [Li21] have confirmed Nishikawa’s conjecture for the positive and nonnegative cases, respectively. For further information and recent developments on this topic, we direct readers to the following references:[DF24, DFY24, FL24, Li22Kahler, Li22PAMS, Li22product, Li24, NPW22, NPWW22].
@@
Any closed Einstein manifold with nonnegative curvature operator of the second kind is either flat or a round sphere.
There is a rich literature on Theorem 1.1. Kashiwada [Kashiwada93] proved that compact manifold $M$ with harmonic curvature and positive curvature operator of the second kind is a space of constant curvature. Cao-Gursky-Tran [CGT23] demonstrated that compact Einstein manifolds with 4-positive and 4-nonnegative curvature operator of the second kind are constant curvature spaces and locally symmetric spaces, respectively. Subsequently, Li [Li22JGA] relaxed the assumption to $4\frac{1}{2}$-positive (respectively, nonnegative) curvature operator of the second kind. Nienhaus-Petersen-Wink [NPW22] showed that $n$-dimensional compact Einstein manifolds with $N^{\prime}$-nonnegative ($N^{\prime}<\frac{3n(n+2)}{2(n+4)}$) curvature operators of the second kind are either rational homology spheres or flat. Recently, Dai, Fu and Yang [DFY24] showed that a compact Einstein manifold of dimension $n\geq 4$ with nonnegative curvature operator of the second kind is flat or a space form of positive curvature, see also [DF24].
Cao-Gursky-Tran [CGT23], Nienhaus-Petersen-Wink [NPW22], Li [Li21] and Dai-Fu [DF24] have proved many excellent results under various curvature conditions. Recently, Li [Li24] introduced new lower bound conditions on the curvature operator of the second kind (see [Li24, Definition 1.2]) and proved optimal sphere theorems. Based on this framework, we focus on Li’s condition with $\alpha=1$, namely (1.1), which is weaker than nonnegative curvature operator of the second kind. To state our result, we introduce some notation: $(M^{n},g)$ is an $n$-dimensional Riemannian manifold, $\mathring{R}$ represents the curvature operator of the second kind, N=(n+2)(n1)/2N=(n+2)(n-1)/2 is the dimension of $S_{0}^{2}(T_{p}M)$, $\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{N}$ are the eigenvalues of $\mathring{R}$, and $\bar{\lambda}=\sum_{j=1}^{N}\lambda_{j}/N$ is the average of $\lambda_{j}$. The main result of this paper is as follows:
@@
(1.1) $\lambda_{1}\geq-\theta(n)\bar{\lambda},$
@@
(1.2) $\theta(n)=\begin{cases}17/7,&\quad n=4,\\ 47/18,&\quad n=5,\\ \frac{2(n^{4}-4n^{3}-5n^{2}+30n-16)}{n^{4}+8n^{3}-5n^{2}-48n+32},&\quad n\geq 6.\end{cases}$
@@
The crucial step in proving Theorem 1.2 is solving an extremum problem introduced in Proposition 3.4. Upon examination of the proof, it becomes evident that for any $\theta$, strictly greater than $\theta(n)$, Proposition 3.4 does not hold. Consequently, by employing our method, $\theta(n)$ as provided in equation (1.2) is deemed optimal. But it remains an interesting question that whether $\theta(n)$ is sharp or not for Theorem 1.2. Recently, the authors [CW25] obtained analogous results under a cone condition where the first two eigenvalues have a negative lower bound.
@@ 2.1. Notation and Conventions
In this section, we set up some notations and recall some basic facts about the curvature operator of the second kind, and refer to [BK78, CV60, Kashiwada93, Li21, Nishikawa86, OT79] for more details.
@@ 2.2. Curvature Operator of the Second Kind
For an algebraic curvature tensor $\in S^{2}_{B}(\wedge^{2}(T_{p}M)), RR$ can induce two self-adjoint operators (see [Nishikawa86]). The first one denoted by $\hat{R}:\wedge^{2}(T_{p}M)\to\wedge^{2}(T_{p}M)$ is defined as
@@
The curvature operator of the second kind, introduced by Nishikawa [Nishikawa86], is symmetric bilinear form
@@
obtained by restricting $\overline{R}$ to $S^{2}_{0}(T_{p}M)$, the space of traceless symmetric two-tensors. It was pointed out in [NPW22] that the curvature operator of the second kind $\mathring{R}$ can also be interpreted as the self-adjoint operator
@@
Suppose $(M^{n},g)$ is an Einstein manifold of dimension $n$. Then the scalar curvature satisfies
@@ 2.3. A formula for Weyl tensor on Einstein manifolds
[…]
$(A\mathchoice{\mathbin{\vtop{\halign{#\cr$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}}{}}B)_{ijkl}=A_{ik}B_{jl}+A_{jl}B_{ik}-A_{jk}B_{il}-A_{il}B_{jk}.$
It is well known that the Riemann curvature tensor $R$ can be decomposed into irreducible components (cf. [CaM20, (1.79)]) as
(2.19) $R=W+\frac{1}{n-2}\operatorname{Ric}\mathchoice{\mathbin{\vtop{\halign{#\cr$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}}{}}g-\frac{{\operatorname{Scal}}}{2(n-1)(n-2)}g\mathchoice{\mathbin{\vtop{\halign{#\cr$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}}{}}g,$
@@
(2.28) $R=W+\frac{{\operatorname{Scal}}}{2n(n-1)}g\mathchoice{\mathbin{\vtop{\halign{#\cr$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}}{}}g.$
@@
The proof of Theorem 1.2 relies on establishing the key estimate
$\langle\Delta R,R\rangle\geq 0,$
where $R$ denotes the Riemann curvature tensor. Our approach begins by expressing $\langle\Delta R,R\rangle$ in terms of the curvature operator of the second kind under the cone condition (1.1) for Einstein manifolds.
We recall some curvature notations on $(M^{n},g)$: $R$ represents the Riemann curvature operator, $W$ represents the Weyl tensor, $\mathring{R}$ represents the curvature operator of the second kind, $\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{N}$ are the eigenvalues of $\mathring{R}$, and $\bar{\lambda}=\sum_{j=1}^{N}\lambda_{j}/N$ is the average of $\lambda_{j}$. To prove the main theorem, we conduct several computations involving the Riemann curvature through the curvature operator of the second kind on Einstein manifolds.
@@
Let $(M^{n},g)$ be an $n$-dimensional Einstein manifold. Then
@@
$R=\frac{{\operatorname{Scal}}}{2n(n-1)}g\mathchoice{\mathbin{\vtop{\halign{#\cr$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}}{}}g+W,$
@@
$|R|^{2}=\frac{{\operatorname{Scal}}^{2}}{4n^{2}(n-1)^{2}}|g\mathchoice{\mathbin{\vtop{\halign{#\cr$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}}{}}g|^{2}+|W|^{2}=\frac{{\operatorname{Scal}}^{2}}{2n(n-1)}+|W|^{2},$
[…]
@@
Let $(M^{n},g)$ be an Einstein manifold of dimension $n$. Suppose the smallest eigenvalue of $\mathring{R}$ satisfies $\lambda_{1}\geq-\theta\bar{\lambda}$, then
@@
and by assumption $\lambda_{1}\geq-\theta\bar{\lambda}$, we estimate that
@@
Let $(M^{n},g)$ be an $n$-dimensional Einstein manifold. Suppose the curvature operator of the second kind satisfies $\lambda_{1}\geq-\theta\bar{\lambda}$. Then
@@
$=\quad$ $\frac{16}{3n}\Big[N(N-3)\theta-(2N-9n+6)N\Big]\bar{\lambda}^{3}$
@@
Let $(M^{n},g)$ be an $n$-dimensional Einstein manifold. Suppose the curvature operator of the second kind satisfies (1.1). Then
(3.6) $\left\langle\Delta R,R\right\rangle\geq 0,$
with the equality holding if and only if
$\lambda=(1,1,\cdots,1)\bar{\lambda}\text{\qquad or \qquad}\lambda=(-\theta,\frac{N+\theta}{N-1},\cdots,\frac{N+\theta}{N-1})\bar{\lambda}.$
@@
By the curvature condition (1.2), we have $\bar{\lambda}\geq 0$. If $\bar{\lambda}=0$, then $\lambda_{1}=\lambda_{2}=\cdots=\lambda_{N}=0$, and (3.6) holds trivially by (3.5). Thus, we focus on the case $\bar{\lambda}>0$.
@@
We prove the case of $n\geq 6$. From the expression of $f(\lambda)$ we can deduce that
@@
$\sum_{j=1}^{N}x_{j}=Nt,\quad\text{and}\quad x_{j}\geq 0\ (j=1,2,\cdots,N).$
@@
When $l$ equals $k/2$, we observe that $1-\frac{N-2}{3(N-1)}<\frac{N}{k}$, and in this case equation (3.8) becomes unsolvable. When $l<\frac{k}{2}$, by solving (3.8) we get
@@
Therefore we conclude from (3.9) that
$\min\left\{F(x):\sum_{j=1}^{N}x_{j}=Nt,x_{j}\geq 0\right\}=\min\left\{F(P_{k,0}),k=1,2,\cdots,N\right\},$
@@
$\lambda^{m}=\left(\underbrace{-\theta\bar{\lambda},\cdots,-\theta\bar{\lambda}}_{m},\underbrace{\frac{N+m\theta}{N-m}\bar{\lambda},\cdots,\frac{N+m\theta}{N-m}\bar{\lambda}}_{N-m}\right)$
@@
Recall from the definition of $\theta$ in (1.2) that
$\theta=\frac{2\left(n^{4}-4n^{3}-5n^{2}+30n-16\right)}{n^{4}+8n^{3}-5n^{2}-48n+32}=\frac{3(N-1)(N+1-n)}{(N-1)(N-3)+3n(N-2)}-1,$
we see
@@
and substituting above to (3.10) yields
$f(\lambda^{0})=f(\lambda^{1})=0,\quad\text{and}\quad f(\lambda^{k})>0$
for $k=2,3,\cdots,N-1$. Hence Proposition 3.4 comes true for $n\geq 6$.
For dimensions $n=4$ and $5$, $\langle\Delta R,R\rangle$ can be expressed explicitly in terms of the curvature operator of the second kind as (see [DF24, Section 4.2]):
(3.11) $\left\langle\Delta R,R\right\rangle=\begin{cases}8\left(\sum_{j=1}^{9}\lambda_{j}^{3}-9\bar{\lambda}^{3}\right),&\quad n=4,\\ 8\left(\sum_{j=1}^{14}\lambda_{j}^{3}+\frac{1}{3}\bar{\lambda}\sum_{j=1}^{14}\lambda_{j}^{2}-\frac{56}{3}\bar{\lambda}^{3}\right),&\quad n=5.\end{cases}$
Then an argument similar to the one used for $n\geq 6$ shows that Proposition 3.4 holds for $n=4$ and $n=5$. This completes the proof of Proposition 3.4.
$\square$
Now we prove the main theorem.
Proof of Theorem 1.2.
Using (3.6), we have
@@
and integrating the above inequality over $M$ yields
$0=\int_{M}\Delta|R|^{2}\,d\mu_{g}\geq\int_{M}2|\nabla R|^{2}\,d\mu_{g}\geq 0,$
which implies $|\nabla R|=0$. Therefore $M$ is a symmetric space. [NPW22, Theorem B] and [Li24, Theorem 1.8(2)] imply that $M$ is either flat or a rational homology sphere. Besides, Poincaré duality shows that rational cohomology is determined by rational homology [Hatcher02, Chapter 3.3]. Thus we conclude that $M$ is either flat or a symmetric space which is also a rational cohomology sphere. Compact symmetric spaces which are real cohomology spheres were classified by Wolf in [Wolf, Theorem 1]. If $\bar{\lambda}>0$, apart from spheres, $SU(3)/SO(3)$ is the only simply connected example. Proposition 3.4 implies that the eigenvalues of $\mathring{R}$ is either $\lambda=(1,\cdots,1)\bar{\lambda}$ or $\lambda=(-\theta,\frac{N+\theta}{N-1},\cdots,\frac{N+\theta}{N-1})\bar{\lambda}$, and combining with [NPW22, Example 4.5], we know that $M$ is a round sphere. $\square$